OpenStudy (anonymous):

Two factory plants are making TV panels. Yesterday, Plant A produced 3 times as many panels as Plant B. Three percent of the panels from Plant A and 4% of the panels from Plant B were defective. How many panels did Plant A produce, if the two plants together produced 1300 defective panels?

OpenStudy (anonymous):

Known equations: A=3B .03A +.04B = 1300 Substitute A=3B into second equation: .03(3B) + .04B = 1300 .09B + .04B = 1300 .13B = 1300 B = 1300/.13 = 10,000 panels A = 3(10,000) = 30,000 panels **Plant A produced 30,000 panels Double check: .03(30,000) = 900 .04(10,000) = 400 900+400=1300 defective panels

OpenStudy (anonymous):

Is there any other method besides the substitution method to do this? In class I don't think we did the substitution method. Thank you soooo much though.

OpenStudy (anonymous):

Substitution would be the easiest way, since you already have an equation where a one variable equals something (A=3B) You could try doing it by elimination, but in my opinion that's much more tedious and harder to do since you're dealing with decimals.

OpenStudy (anonymous):

Is there a way I can get rid of decimals? Like, wouldn't I multiply everything by 100 in order to make it a whole number?

OpenStudy (anonymous):

Elimination: A-3B = 0 .03A + .04B = 1300 Multiply second equation by 100: A-3B = 0 3A +4B = 130000 Easiest way is to multiply top equation by -3 -3A+9B = 0 3A+4B = 130000 Add: 13B = 130000 B = 10,000 Then just do the same steps above to find A again.

OpenStudy (anonymous):

Ok, that helps a lot. Thank you so much c: